OpenStudy (anonymous):

Find all solutions to the equation. 7 sin2x - 14 sin x + 2 = -5

OpenStudy (anonymous):

secret letter E (EVERYTHING)

OpenStudy (campbell_st):

this is a quadratic equation... start by letting it equal zero \[7(\sin^2(x) -2\sin(x) + 1) = 0\] you'll find the quadratic is a perfect square

OpenStudy (anonymous):

I got that but at this point I am lost. Unless that is the answer which I don't believe it is.

OpenStudy (campbell_st):

ok... so it factors to \[7(\sin(x) -1)^2 = 0\] so what values of x make sin(x) - 1 = 0

OpenStudy (anonymous):

X=1

OpenStudy (campbell_st):

nope sin(x) = 1 \[x = \sin^{-1} (1)\] whats the angle

OpenStudy (anonymous):

1.5707

OpenStudy (campbell_st):

ummm are you working in degrees or radians...?

OpenStudy (anonymous):

degrees

OpenStudy (campbell_st):

ok... so the only place when sin(x) = 1 is where x = 90 degrees

OpenStudy (anonymous):

ok